Differences between revisions 8 and 9
Revision 8 as of 2022-06-24 14:26:14
Size: 2741
Revision 9 as of 2022-06-24 14:26:56
Size: 2740
Deletions are marked like this. Additions are marked like this.
Line 45: Line 45:
 * Monitor of the build up. Height of the top chamber was changed by 1mm.  * Monitored the build up. Height of the top chamber was changed by 1mm.


Zoom Meeting

CRY+VIS Interface Meeting

Date: 2022/6/24 13:00-13:50

Participants: Takahashi, Sato, Uchiyama, Yokozawa, Aso, Kimura, Miyoki, Ushiba, Akutsu, Ikeda

Post O3 Upgrade portal


  • IX: Heater control of the laser for OpLev.

  • EX:
  • IY: Cooling of the ducts is ready.
  • EY: Vent and monitor. Repair of leak and pump-down in the next week.
  • Type-B: Stand-alone SM driver for SR3. FC to SR3/SRM. Investigation of the mirror replacement in SRM. Installation of the MSB to SR3/SRM in the next week.
  • Type-Bp: Test of the traverser limit SW in PR2/PR3. Health check. OSEM and OpLev adjustment in PRM.

  • Type-C: Health check should be scheduled.


  • Replacement of flip cables: EXT and BF temp.
  • TM height setting:
  • Oplev: Investigating the relation between laser and temperature. Adjustment for PRM. For SRs in the next week.
  • ACC:
  • Commissioning in air:
  • Commissioning in vacuum:
  • Acceptance check:
  • PF actuator: At the in-vacuum commissioning.
  • New vacuum pumps for V-chambers:
  • Heater control: Pending.
  • TMS: Window to EYT in the next week.


VIS Meeting

Date: 2022/6/24 14:00-14:15

Participants: Takahashi, Uchiyama, Ushiba, Fabian, Tanaka, Aso, Washimi

Summary Report


  • Monitored the build up. Height of the top chamber was changed by 1mm.


  • We inspected SRM payload and chamber and we think we can exchange the small mirror after the installation of the mid-size baffle. When are we going to change it?

  • In SR3, we locked the suspension and First Contact was applied.
  • In the corresponding filter bank, I updated the information of the transfer function of the new geophone.


  • In PR3 and PR2, we tested the Traverser limit switches.
  • In PRM, OSEMs H2 and H3 were close to the end of their linear ranges, therefore, we moved them. The payload is not as stable as we would've liked, and was disturbed when we installed the Traverser limit switches. See klog.
  • Health check of PRs.





  • Type-A: Monitor of pump down.
  • Type-B: MSB in SR3/SRM. OpLev setting.

  • Type-Bp: Health check.
  • Type-C:



Travel Plans

(Date in Kamioka)

The week from 6/27

  • Takahashi: 6/27-7/1
  • Sato: 6/27-7/1
  • Hirata: 6/28-7/1

The week from 7/4

  • Takahashi: 7/5-8
  • Sato:
  • Hirata:

Next meeting

  • 2022/7/1(FRI)

KAGRA/Subgroups/VIS/VISMinutes20220624 (last edited 2022-06-24 14:26:56 by RyutaroTakahashi)